OpenStudy (anonymous):

The radius of a circle is multiplied by 2. Its area is _____ times bigger than the original circle. 3 4 6 9

OpenStudy (anonymous):

The formula is A = \(\pi\)r², right? If you twice the radius, then its area is \(\pi\)(2r)² = 4\(\pi\)r². Original circle: \(\pi\)r² New circle: 4\(\pi\)r² So the answer is 4.

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me with more please

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

@tafkas77 If you assume Rebecca no longer need your help, just go and help other users. There always will be other users there for her to help her. You don't have to worry about her all the time. Thanks.

OpenStudy (anonymous):

Thanks @RD⁴² I get your point, but I take this rather seriously sometimes. Rebecca constantly asks for my help, so I try to help her whenever I can, because I know she's come to depend on me for math help. Your advice is good and warranted, however, I thought she should know. Many people (not just me) get frustrated with this, so I thought I'd let her know. Thanks! :)